OpenStudy (anonymous):

Help Pwease! ^^ What are all the multiplication pairs for 27? ((I'm not a cheating 5th grader... I'm currently answering a trinomial factoring question... I'm lazy *sigh*.)) Thanks!

jimthompson5910 (jim_thompson5910):

you mean all the ways to factor 27?

OpenStudy (anonymous):

1*27, 9*3

OpenStudy (anonymous):

Yes as in prime factorization.

jimthompson5910 (jim_thompson5910):

if so, then 27 = 1*27 27 = 3*9 you can reorder the terms however you want since order doesn't matter

OpenStudy (anonymous):

The prime factorization would be 3*3*3

OpenStudy (anonymous):

Ah, okay thanks both!

jimthompson5910 (jim_thompson5910):

and you can write 3*3*3 as 3^3, either one works

OpenStudy (anonymous):

Okay, thank you so much!

jimthompson5910 (jim_thompson5910):

yw